A buffer solution with a volume of 0.0250 L consists of 0.61 M iodous acid (HIO2), a weak acid, plus 0.61 M lithium iodite (LiIO2). The acid dissociation constant of iodous acid, Ka, is 3.2 ✕ 10−5. Determine the pH of the buffer solution after the addition of 0.0034 mol sodium hydroxide (NaOH), a strong base. (Assume no change in solution volume.)

can u solve with an ice table? thank you.
Answer #1

Answer:

To determine pH

Consider,

Moles of HIO2 = 0.61*0.0250

= 0.01525

HIO2 left after adding KOH = 0.01525 - 0.0034

= 0.01185 mol

IO2- = 0.01525 + 0.0034

= 0.01865

Now consider,

3.2*10^-5 = 0.01525*[H3O+] / 0.01185

[H3O+] = 3.9*10^-5

pH = -log(3.9*10^-5)

pH = 4.41